the sin θ = -3/4, tan θ is positive. find the other five functions ( exact values)
Answer by Alan3354(69443)

the sin θ = -3/4, tan θ is positive. find the other five functions ( exact values)
-----------
The sine is negative in Quads 3 & 4
The tan is positive in Quads 1 & 3
--> the angle is in Quad 3
-----------------
The hypotenuse = 4, x = -3, y = - sqrt(7)
cos = -3/-sqrt(7) = 3sqrt(7)/7
tan = sin/cos = (-3/4)/(3sqrt(7)/7 = -3sqrt(7)/-21 = sqrt(7)/7
-------------
You can do the remaining 3
cot = 1/tan
sec = 1/cos
csc = 1/sin